OpenStudy (anonymous):

find the limit as x approaches -6 of the function (2x+12)/|x+6| if it exists. If the limit does not exist, explain why.

OpenStudy (dumbcow):

lim = 0/0 differentiate top and bottom derivative of top = 2 since bottom is absolute value it cannot be differentiated at x=-6 so we split it into 2 cases where x<-6 and x>-6 x<-6: derivative = -1 x>-6: derivative= 1 \[\lim_{x \rightarrow -6^{-}} = \frac{2}{-1}=-2\] \[\lim_{x \rightarrow -6^{+}} = \frac{2}{1}= 2\] Because the 2 limits do not agree the lim x->-6 does not exist